OpenStudy (anonymous):

y = 5x and 2x + 3y = 34

OpenStudy (anonymous):

Choose the correct first step(s) to solve the system using substitution.

OpenStudy (anonymous):

2x+3(5x)=34 17x=34 x=2 y=5(2)=10

OpenStudy (mimi_x3):

Sub it in: \[2x+3(5x)=34\] \[2x+15x=34\] \[17x=34\] \[x = 34/17\] =>\(x=2\)

OpenStudy (anonymous):

that is correct! Thank you